# literature.bib

### v2

@@ 5.2.2 Image Denoising
We now sample from the posterior of an image denoising task. Due to the higher dimension, we can not compute accurate approximations of the Wasserstein distances to the target anymore. We therefore resort to illustrating the behaviour of ULPDA using moment estimates. A ground truth image normalized to intensities $[0,1]$ is distorted by noise with standard deviation $0.25$ (PSNR 12.0212.0212.02\,12.02dB in the noisy observation). We then draw $10^{6}$ samples with a single chain and compute the sample mean and pixel-wise variance after a burn-in phase of $31e$ 3 samples. The experiment is repeated three times for values $\lambda=1,10,100$, each with the same step size $\tau$. A fourth chain is computed using Prox-Sub with the same $\tau$. MMSE and pixel-wise variance estimates are reported in Figure 5. The sample mean (MMSE) estimates produced by the different chains converge quickly and are very similar for all chains (PSNR values between 23.8223.8223.82\,23.82dB and 23.8523.8523.85\,23.85dB). The pixel-wise variance, however, shows the same effect of overdispersion for smaller levels of $\lambda$ as in the lower-dimensional example.
While the ‘primal’ samples $X^{n}$ are overdispersed for finite $\lambda$, the dual iterates $Y^{n}$ characteristically exhibit underdispersion (see Figure 6). This effect was consistent throughout all our numerical experiments (compare also Figures 2 and 3). A heuristic explanation for overdispersion in the primal and underdispersion in the dual variable lies in the fact that the step in the dual variable solves the optimality condition $y\in\partial f(Kx)$ only incompletely. On the particle level in continuous time, when the primal variable $X_{t}$ is driven towards the tail regions of its marginal distribution by the Brownian motion, the dual variable is not evolved entirely to the corresponding position $\partial f(KX_{t})$ in the tails of its marginal, but rather remains near the mode of its marginal. The effect decreased with growing parameters $\lambda$ in all our experiments.
@@ 5.3 Total Generalized Variation Based Image Denoising
In this experiment, we consider an example with an image prior based on second order total generalized variation (TGV). Like the TV functional, total generalized variation allows for discontinuous edges in images, but it generalizes TV in that it also employs higher-order derivatives. This way, it avoids the typical staircasing effect of TV in smooth but non-constant regions of images. For optimization purposes, the second order TGV functional is usually defined as
@@ 6 Conclusion and Open Problems
We have analyzed a recently proposed algorithm for sampling from log-concave, non-smooth potentials which makes use of alternating primal-dual steps as is well-known in convex non-smooth optimization. We showed that the sampling algorithm corresponds to a stochastic differential equation and formulated the mean field limit Fokker-Planck equation. Using a simple coupling argument, contraction of the continuous-time limit to a stationary solution was proved. Unlike for purely primal overdamped Langevin diffusion, the primal-dual Fokker-Planck equation does not have the target distribution as its stationary solution. A proposed correction of the Fokker-Planck equation forcing the target to be its stationary solution is possible, but eventually not useful for numerical implementation. However, the target distribution can be related to the stationary solution of the uncorrected scheme in the limit of the ratio of primal and dual step sizes growing to infinity. We made use of coupling techniques to prove a corresponding bound in Wasserstein distance to the target for a finite step size ratio. With similar coupling arguments in discrete time, estimates in the typical spirit of convergence analysis of the optimization algorithm can be transferred to the sampling case, showing stability of the sampling method in Wasserstein distance.
@@
We are currently working on a final coupling result, filling the gap between the stationary solution in discrete time and the stationary solution in continuous time. This, together with the stability in discrete time and the bound between continuous-time stationary state and the target would allow a convergence analysis between the algorithm and the target. Numerically, the results suggest that if there is a way to compute the subgradient of the dualized potential term, the additional bias due to the finite step size ratio makes the primal-dual samples consistently over-dispersed compared to the method of [Habring2023]. We currently see the primal-dual method’s use cases in two settings. If the subgradient of the dualized term is not available, but the proximal mapping of its convex conjugate is, the method is up to our knowledge the first applicable algorithm able to draw samples from an approximation of the posterior. Secondly, if the additional bias due to a larger finite step size ratio does not matter (because a correction step is added or because downstream tasks can be carried out with biased samples), the method can be applied with potentially larger step sizes than a purely primal method, allowing for faster mixing time. To find out whether this makes the primal-dual method more efficient in applications, further numerical studies, in particular in combination with a Metropolis-Hastings correction step, will be needed.

### v3

@@ 5.2.2 Image Denoising
We now sample from the posterior of an image denoising task. Due to the higher dimension, we can not compute accurate approximations of the Wasserstein distances to the target anymore. We therefore resort to illustrating the behaviour of the algorithm (1.4) using moment estimates. A ground truth image normalized to intensities $[0,1]$ is distorted by noise with standard deviation $0.25$ (PSNR 12.0212.0212.02\,12.02dB in the noisy observation). We then draw $10^{6}$ samples with a single chain and compute the sample mean and pixel-wise variance after a burn-in phase of $31e$ 3 samples. The experiment is repeated three times for values $\lambda=1,10,100$, each with the same step size $\tau$. A fourth chain is computed using Prox-Sub with the same $\tau$. MMSE and pixel-wise variance estimates are reported in Figure 6. The sample mean (MMSE) estimates produced by the different chains converge quickly and are very similar for all chains (PSNR values between 23.8223.8223.82\,23.82dB and 23.8523.8523.85\,23.85dB). The pixel-wise variance, however, shows the same effect of overdispersion for smaller levels of $\lambda$ as in the lower-dimensional example.
While the ‘primal’ samples $X^{n}$ are overdispersed for finite $\lambda$, the dual iterates $Y^{n}$ characteristically exhibit underdispersion (see Figure 7). This effect was consistent throughout all our numerical experiments (compare also Figures 3 and 4). A heuristic explanation for overdispersion in the primal and underdispersion in the dual variable lies in the fact that the step in the dual variable solves the optimality condition $y\in\partial f(Kx)$ only incompletely. On the particle level in continuous time, when the primal variable $X_{t}$ is driven towards the tail regions of its marginal distribution by the Brownian motion, the dual variable is not evolved entirely to the corresponding position $\partial f(KX_{t})$ in the tails of its marginal, but rather remains near the mode of its marginal. The effect decreased with growing parameters $\lambda$ in all our experiments.
@@ 5.3 Total Generalized Variation Based Image Denoising
In this experiment, we consider an example with an image prior based on second order total generalized variation (TGV). Like the TV functional, total generalized variation allows for discontinuous edges in images, but it generalizes TV in that it also employs higher-order derivatives. Acting as a penalty also on higher order image gradients, TGV-based MAP estimates avoid the typical staircasing effects of TV in smooth but non-constant regions of images. For optimization purposes, the second order TGV functional is usually defined as
@@ 6 Conclusion and Open Problems
We have analyzed a recently proposed algorithm for sampling from log-concave, non-smooth potentials which makes use of alternating primal-dual steps as is well-known in convex non-smooth optimization. We showed that the sampling algorithm corresponds to a stochastic differential equation and formulated the mean field limit Fokker-Planck equation. Using a simple coupling argument, contraction of the continuous-time limit to a stationary solution was proved. Unlike standard overdamped Langevin diffusion, the primal-dual Fokker-Planck equation is not a gradient flow in the corresponding Wasserstein space and does not have the target distribution as its stationary solution. A proposed correction of the Fokker-Planck equation forcing the target to be its stationary solution is possible, but eventually not useful for numerical implementation. However, the target distribution can be related to the stationary solution of the uncorrected scheme in the limit of the ratio of primal and dual step sizes growing to infinity. For the unmodified scheme, we made use of coupling techniques to prove a bound on Wasserstein distance between the stationary distribution and the target for a finite step size ratio.
With similar coupling arguments in discrete time, estimates in the typical spirit of convergence analysis of the optimization algorithm can be transferred to the sampling case. This allowed us to show stability of the sampling method in Wasserstein distance, resulting in the existence of a stationary measure in discrete time under strong convexity assumptions. Under the additional assumption of Lipschitz-smoothness of one of the potentials, we can bound the discretization error using a coupling between discrete and continuous time. This final bound implies a convergence result that relates the generated samples with the target. An important open question is whether the regularity assumptions on the potentials $f,g$ in the stability result and the discretization error bounds can be relaxed to not require differentiability.
@@
Numerically, the results suggest that if there is a way to compute the subgradient of the dualized potential term, the additional bias due to the finite step size ratio makes the primal-dual samples consistently over-dispersed compared to discretizations of overdamped Langevin diffusion like the method of [Habring2023]. We currently see the primal-dual method’s use cases in two settings. If the subgradient of the dualized term is not available, but the proximal mapping of its convex conjugate is, the method is up to our knowledge the first applicable algorithm able to draw samples from an approximation of the posterior. Secondly, if the additional bias due to a larger finite step size ratio does not matter (because a correction step is added or because downstream tasks can be carried out with biased samples), the method can be applied with larger step sizes than a purely primal method, allowing for faster mixing time. To find out whether this makes the primal-dual method more efficient in applications, further numerical studies. Particularly, the combination of a proposal step using the primal-dual algorithm and a Metropolis-Hastings correction step seems like a promising method and should be analyzed further in future work.